Twenty years ago, 57% of parents of children in high schule" was a serious probem that high schoo s udens were not being taugt enagh man and science. A recent survey found that 297 of 750 parents of children in high school felt it was a serkous problem that high school students were not being taught enough math and science. Do parents feel difterenty hoday than they ded twenty years ago? s the r 0,0 level of significance. Because npu (1-po)= 10, the sample size is 5% of the population size, and the sample the requirements for testing the hypothesissatistied Round to one decimal placI ctv the answer box and then click Check Answer Enter your answer in Check Anawer parts Clear A
s y Twenty years ago, 57% of parents of children in high school fet it was a serious problem that high school students were not being and science. A recent survey found that 297 of 750 parents of chikdren in being taught envough math and science. Do parents feel differently today than they did twenty years ago?Ue th 005 level of taught enough math high school felt it was a serious problem that high school students were not significance. Becausen@po-popu-, 10the sar piesze is!- 5%ofthe populatkin size, andthe sample he hypothesis the ▼: satisfied Round to one decimal place as needed.) ess than greater than Entor your answer in the answer box and then click Check Answer Activa Check An 4
Twenty years ago. 57% of parents of children scilence. A recent survey found that 297 of 750 parents of children in high school felt it was a serkous problem that high school students were not being taught enough math and science. Do parents feel oitherenty koday thian they did wenty years apo? Ube te e-005 level of skgnificance in high school tett it was a serious protiem that h h school students were not being tight enough math ? Use the 0.05 level of про(1-01-110, the sample size is 5% of the population size, and the sample the requirements for testing the hypothesis salisfed is given to be random s given to not be random, can be reasonably assumed to be random, cannot be reasonably assumed to be random, Enter your answer in the answer box and then click Check Answer Activ Check Anaw 4 Cear Al 1 8
Twenty years ago, 57% of parents of children in high stolen it was a senus p enthal high school sto ents were not beng tau tenugh man and science. A recent survey found that 297 of 750 parents of children in high school felt it was a serkous problem that high school students were not being taught enough math and science. Do parents feel İff rent y today than they do twenty years ago? Use he a 0 05 level of significance. Because nPo(1-por-E10, the sample size is L.! "İ 5% of the population size, and the sample tihe requirementsfor testing the hypotes-dsa sted. Round to one decimai place as needed.) are not Activa Enter your answer in the answer box and then click Check Answer Check Answer lear Al 4 romaining
Answer #1

np(1-p) = 750*0.57*(1-0.57)=183.8

because np(1-p) = 183.8 > 10, the sample size is less than 5% of population size, and sample can be reasonably assumed to be random , the rquirement for testing the hypothesis is satisfied